# Practical security of twin-field quantum key distribution under wavelength-switching attack

### v1

@@ III.2 Testing of wavelength-switching attack
In a TF-class QKD system, Charlie provides the reference light to Alice and Bob and performs Bell-state measurements on the quantum signals they send back, as depicted by the blue area in Fig. 1(a). Ensuring stable interference is crucial for successful Bell-state measurement in order to achieve high visibility of quantum state detection. This step plays a critical role in establishing secure key generation between Alice and Bob within the TF-class QKD system. Therefore, it is essential to verify the stability of interference under wavelength-switching attacks. To maintain stable interference despite such attacks, Eve manipulates the wavelength of the reference laser while simultaneously affecting Alice’s and Bob’s lasers. The key lies in selecting an appropriate switching rate ($r_{s}$) that achieves stable interference. Figure 1(b) illustrates Eve’s selection of four switching rates ($r_{s}$) as attacking parameters. The black and red solid lines represent optical power measured by PD3 and PD4 at two output arms of Charlie’s interference setup respectively. When there is no attack ($r_{s}=0$), Charlie measures smooth and stable interference over time. However, slight fluctuations are observed when Alice’s and Bob’s wavelengths are switched at rates of either $100$ $\mathrm{k}\mathrm{Hz}$ or $200$ $\mathrm{k}\mathrm{Hz}$, indicating potential disturbance caused by these attacks.
@@
III.3 Gain of average photon number under attack
@@
Figure 4: The experimental results demonstrated the average photon number gain of the TF-class QKD system in the presence of a wavelength-switching attack.
The correlation between the shifted wavelength under attack and the gain in mean photon number of emitted quantum states in the TF-class QKD system is illustrated in Fig. 4. The experimentally measured values after conversion are denoted by yellow markers. The gain in mean photon number of the TF-class QKD system exhibits a rapid increase when $f_{\Delta}$ is less than or equal to $10~{}\mathrm{M}\mathrm{Hz}$. Beyond this threshold, the rate of increase slows down. Therefore, we limit our frequency shift to a maximum of $30~{}\mathrm{M}\mathrm{Hz}$ during experiment, resulting in an 8.7% enhancement of mean photon number in the TF-class QKD system.
IV Effect on the security of TF-class QKD
@@
IV.1 Security analysis on TF QKD
In the TF QKD system, Alice and Bob generate phase-randomized optical fields, which are transmitted to Charlie through a quantum channel. Subsequently, Charlie performs the Bell-state measurement and announces an event only if a single-photon detector is clicked. The event will be considered valid for generating a secure key if Alice and Bob select the same random phase. To evaluate the potential vulnerability of wavelength-switching attacks, we consider typical implementations where Alice and Bob employ three different intensities: $\mu_{0}$, $\mu_{1}$, and $\mu_{2}$. Here, $\mu_{0}$ represents the signal state while $\mu_{1}$ and $\mu_{2}$ represent decoy states. In scenarios with a realistic number of transmitted signals approaching asymptotic limits, it is possible to establish a lower bound on the secret key rate [6].
$R_{L}=\frac{d}{M}\left\{Q_{1,L}^{\mu_{0}}[1-H_{2}(e_{1,U})]-fQ_{u}H_{2}(E_{u})\right\},$ (1)
where $d$ is the duty cycle between the classical and the quantum modalities, $M$ is the value of phase slices, $Q_{1,L}^{\mu_{0}}$ denotes a lower bound on the single-photon yield, $e_{1,U}$ denotes an upper bound on the phase error rate, $f$ is the efficiency of error correction, $Q_{U}$ represents the overall experimentally observed gain, $E_{u}$ is the overall experimentally observed quantum bit error rate (QBER), and $H_{2}(x)=-x\log_{2}(x)-(1-x)\log_{2}(1-x)$ is the binary Shannon entropy function.
Figure 5: The secure key rate is plotted as a function of distance for TF QKD, considering two different values of the multiplicative factor: $g=1$ and $g=1.087$. The simulation parameters used in the calculations are as follows: $P_{d}=2.3\times 10^{-11}$ represents the dark count rate per pulse of the detector; $\alpha=0.16~{}\mathrm{d}\mathrm{B}/\mathrm{k}\mathrm{m}$ denotes the loss constant of low-loss fiber; $\eta_{D}=0.29$ represents the overall efficiency of the detection system; $d=0.5$ indicates the duty cycle between classical and quantum modalities; $M=16$ corresponds to the number of phase slices, and finally, $f=1.16$ signifies the error-correction efficiency.
Figure 6: The secure key rate is plotted as a function of the parameter $f_{\Delta}$ for a fixed distance of $833.8~{}\mathrm{k}\mathrm{m}$ in the TF QKD system. The simulation incorporates the following parameters for calculations. $P_{d}=1.4\times 10^{-11}$, $f=1.16$, $\alpha=0.157$, $\eta_{D}=0.65$, $d=0.5$, and $M=16$.
To estimate $Q_{1,L}$ and $e_{1,U}$, one can use analytical or numerical tools. Here we have that
$Q_{1,L}^{\mu_{0}}\geq\mu_{0}e^{-\mu_{0}}\eta,$ (2)
where $\eta=10^{-\frac{\alpha L/2}{10}}$.
@@ IV.1 Security analysis on TF QKD
In the presence of a wavelength-switching attack, Alice and Bob estimate $e_{1,U}$, $Y_{0,L}$, and $e_{1,U}$ using Eq. 3 and Eq. 4, but now with the experimentally observed quantities $\mu^{\prime}_{1}$, $\mu^{\prime}_{2}$, and $\mu^{\prime}_{3}$. We conduct a simulation analysis using parameters corresponding to a typical experiment [21] to illustrate the security threat posed by a wavelength-switching attack on the TF QKD system. The results with and without an attack on the TF QKD system are depicted in Fig. 5. The black solid line represents the lower bound $R_{L}$ as defined by Eq. 1 under no attack conditions. Subsequently, we simulate the degradation of security bounds due to Eve’s wavelength-switching attack. Specifically, in Fig. 5, the red-dashed and blue-dotted lines respectively indicate values of incorrectly estimated ($R_{L}$) by Alice and Bob when $g=1.087$ in the presence of an attack. As shown in Fig. 5 compared to no-attack scenario represented by black solid line, the secure key rate ($R_{L}$) given by red-dashed line is significantly higher. In other words, Alice and Bob overestimate their secure key rate in response to an attack. The gap between these two curves indicates that Eve can eavesdrop on a portion of secure key rate without being detected while eventually deducing shared secure key between Alice and Bob.
@@
In Fig. 6, we present the impact of the wavelength-switching factor $f_{\Delta}$ on the lower bound of secure key rate at a fixed distance ($833.8~{}\mathrm{k}\mathrm{m}$). The graph illustrates both incorrect and correct estimates of secure key rates for the TF QKD system under attack, represented by red-square and blue-dot points respectively. In this scenario, an increase in parameter $f_{\Delta}$ results in a higher mean photon number for quantum states prepared by Alice and Bob. This consistently elevated mean photon number leads to an overestimation of the lower bound $R_{L}$ of key rate, as indicated by the red-square markers, thereby widening the gap between overestimated and correct key rates illustrated by blue-dot markers. To summarize, wavelength-switching attacks cause Alice and Bob to overestimate their key rate, compromising the security of a TF QKD system.
IV.2 Security analysis on SNS-TF-QKD
@@
Next, we consider the case of SNS-TF-QKD. Similar to the previous example, it is assumed that both Alice and Bob employ three distinct intensities, namely $\mu_{0}$, $\mu_{1}$, and $\mu_{2}$. In a finite-size $N$ of transmitted signals, the lower bound for the secure key rate can be derived from [34].
$R_{L}=$ $2p(1-p)Q_{1,L}^{\mu_{0}}S_{1,L}\left[1-H_{2}(e_{1,U})\right]$ (5)
$-fS_{z}H_{2}(E_{Z}),$
where $p$ is a probability that Alice (Bob) randomly decides to send a phase-randomized coherent state, $S_{z}$ is the counting rate of pulses in $Z$ windows, $S_{1}$ is decoy windows, $E_{z}$ is the corresponding error rate, $H_{2}(x)$, $Q_{1,L}^{\mu_{0}}$, and $e_{1,U}$ follow the previous definition.
To evaluate Eq. 5, Alice and Bob need to calculate the parameters $S_{1,L}$ and $e_{1,U}$ based on the experimentally available data $\mu^{\prime}_{1}$, $\mu^{\prime}_{2}$ and $\mu^{\prime}_{3}$. The lower bound of $S_{1,L}$ is
@@
where $S_{jk}\left(jk=\left\{00,10,01,20,02\right\}\right)$ denotes the counting rate of source $jk$ [34]. The upper bound of the phase-flip error rate $e_{1,U}$ of the untagged photons satisfies [35]
$e_{1,U}\leq\frac{T_{\Delta}-0.5e^{-2\mu_{1}}S_{00}}{2\mu_{2}e^{-2\mu_{1}}S_{1,L}},$ (7)
where $T_{\Delta}$ is the effective rate of detectors.
Figure 7: The secure key rate as a function of distance for SNS-TF-QKD is observed to be $1$ and $1.087$ for two different values of the multiplicative factor $g$. The simulation parameters utilized in the calculations are as follows: $P_{d}=1.4\times 10^{-11}$, $f=1.16$, $\alpha=0.157$, and $\eta_{D}=0.65$.
@@ V Countermeasures
The observed variation in the intensity of quantum states is attributed to the change in optical insertion loss. This change is mainly because the AOM receives an uncharacterized frequency modulation and insufficient driving voltage when the FLC adjusts it to match the reference light’s wavelength switching as analyzed in Sec. III.1. Being immune to the wavelength-switching attack, we propose a calibration procedure for the AOM before system operation to minimize the insertion loss due to driving voltage. In step 1, the driving voltage of the AOM is scanned under a fixed modulation frequency to determine the optimal voltage value, $V_{o}$, corresponding to the minimum insertion loss at this certain modulation frequency. In step 2, the modulation frequency of the AOM is scanned with $V_{o}$ found in Step 1, obtaining the optimal frequency $F_{o}$, corresponding to the further minimized insertion loss. Afterwards, $V_{o}$ and $F_{o}$ are set as the basic parameters for the FLC to control the AOM. This process ensures that the wavelength-switching attack will only increase the insertion loss of the OPLL, thus mitigating the associated security risks.
@@
VI Conclusion
The TF-class QKD has been demonstrated to surpass the fundamental rate-distance limit, which necessitates the generation of identical optical fields by two independent lasers located at separate positions for Alice and Bob. In current implementations, a third-party reference light is commonly employed to lock the wavelengths of the lasers in order to achieve matching. Although this reference-light-locking structure does not directly contribute to key generation, it introduces a potential vulnerability that can be exploited by an eavesdropper, posing a security risk to the devices used by Alice and Bob. This study unveils a wavelength-switching attack on the TF-class QKD system employing an optical OPLL structure. Experimental results demonstrate that Eve can exploit this attack to reduce attenuation in the AOM of the TF-class QKD system, causing optical power from Alice and Bob to exceed intended levels. Importantly, this attack challenges fundamental assumptions in TF-class QKD protocols regarding mean photon number of prepared quantum states and well-protected source stations. Despite altering wavelengths of lights emitted by Alice and Bob, these hacked source units remain consistent, enabling stable interference at Charlie’s location. Theoretical analysis conducted on both TF QKD protocol and SNS-TF-QKD protocol indicates that the wavelength-switching attack could significantly compromise security in implementation of TF-class QKD – legitimate users overestimate secure key rates when facing such attacks. Since this attack is due to imperfect calibration in the external-reference-locking configuration, we propose countermeasures via loss-minimization calibration or using no-external-reference-light scheme [36, 37, 38, 39, 32]. This study demonstrates that the well calibration of TF QKD system is inevitable for its practical security.
Acknowledgements.
We thank Xiangbin Wang and Qiang Zhang for discussions. This work was funded by the Innovation Program for Quantum Science and Technology (2021ZD0300704) and National Natural Science Foundation of China (No. 62371459 and No. 62061136011).
Appendix A Model of AOM

### v2

@@ 2.2.2 Testing of wavelength-switching attack
In a TF-class QKD system, Charlie provides the reference light to Alice and Bob and performs interferometric measurements on the quantum signals they send back, as depicted by the blue area in Fig. 1(a). Ensuring stable interference is crucial for successful interferometric measurement in order to achieve high visibility of quantum state detection. This step plays a critical role in establishing secure key generation between Alice and Bob within the TF-class QKD system. Therefore, it is essential to verify the stability of interference under wavelength-switching attacks. To maintain stable interference despite such attacks, Eve manipulates the wavelength of the reference laser while simultaneously affecting Alice’s and Bob’s lasers. The key lies in selecting an appropriate switching rate ($r_{s}$) that achieves stable interference. Figure 1(b) illustrates Eve’s selection of four switching rates ($r_{s}$) as attacking parameters. The black and red solid lines represent optical power measured by PD3 and PD4 at two output arms of Charlie’s interference setup respectively. When there is no attack ($r_{s}=0$), Charlie measures smooth and stable interference over time. However, slight fluctuations are observed when Alice’s and Bob’s wavelengths are switched at rates of either $100$ $\mathrm{k}\mathrm{Hz}$ or $200$ $\mathrm{k}\mathrm{Hz}$, indicating potential disturbance caused by these attacks.
@@
2.2.3 Gain of average photon number under attack
@@
Figure 4: The experimental results demonstrated the average photon number gain of the OPLL for each source apparatus in the TF-class QKD system under the wavelength-switching attack.
The correlation between the shifted wavelength under attack and the gain in mean photon number of emitted quantum states in the TF-class QKD system is illustrated in Fig. 4. The experimentally measured values after conversion are denoted by yellow markers. The gain in mean photon number of the TF-class QKD system exhibits a rapid increase when $f_{\Delta}$ is less than or equal to 10 MHz. After that, the rate of increase slows down. The gain in mean photon number is maximized at $f_{\Delta}=30~{}\mathrm{M}\mathrm{Hz}$, beyond which the OPLL is no longer working in the locking status. Therefore, we limit our frequency shift to a maximum of $30~{}\mathrm{M}\mathrm{Hz}$ and apply it simultaneously to the OPLLs of both Alice and Bob during the experiment. As a result, the mean photon number of Alice/Bob is increased by 4.35%, and the overall intensity gain is 8.7% for the whole TF QKD system as defined in Ref. [6].
2.3 Effect on the security of TF-class QKD
@@
2.3.1 Security analysis on TF QKD
In the TF QKD system, Alice and Bob generate phase-randomized optical fields, which are transmitted to Charlie through a quantum channel. Subsequently, Charlie performs the interferometric measurement and announces an event only if a single-photon detector is clicked. The event will be considered valid for generating a secure key if Alice and Bob select the same random phase. To evaluate the potential vulnerability of wavelength-switching attacks, we consider typical implementations where Alice/Bob uses three intensity settings $\mu_{a,b}\in\left\{\mu_{0}/2,\mu_{1}/2,\mu_{2}/2\right\}$, with $\mu_{0}=0.4$, $\mu_{1}=10^{-2}$, and $\mu_{2}=10^{-4}$. Here, $\mu_{0}$ represents the signal state while $\mu_{1}$ and $\mu_{2}$ represent decoy states. In scenarios with a realistic number of transmitted signals approaching asymptotic limits, it is possible to establish a lower bound on the secret key rate [6].
$R_{L}=\frac{d}{M}\left\{Q_{1,L}^{\mu_{0}}[1-H_{2}(e_{1,U})]-fQ^{\mu_{0}}H_{2}(E^{\mu_{0}})\right\},$ (1)
where $d=0.5$ is the duty cycle between the classical and the quantum modalities, $M=16$ is the value of phase slices, $Q_{1,L}^{\mu_{0}}$ denotes a lower bound on the single-photon yield, $e_{1,U}$ denotes an upper bound on the phase error rate, $f=1.16$ is the efficiency of error correction, $Q_{U}$ represents the overall experimentally observed gain, $E_{u}$ is the overall experimentally observed quantum bit error rate (QBER), and $H_{2}(x)=-x\log_{2}(x)-(1-x)\log_{2}(1-x)$ is the binary Shannon entropy function.
Figure 5: The secure key rate is plotted as a function of distance for TF QKD, considering two different values of the multiplicative factor: $g=1$ and $g=1.087$.
Figure 6: The secure key rate is plotted as a function of the parameter $f_{\Delta}$ for a fixed distance of $560~{}\mathrm{k}\mathrm{m}$ in the TF QKD system.
To estimate $Q_{1,L}$, $Q^{\mu}$, $e_{1,U}$ and $E_{\mu}$, one can use analytical or numerical tools. Here we have that
$Q_{1,L}^{\mu_{0}}$ $\geq\mu_{0}e^{-\mu_{0}}Y_{1,L},$ (2)
$Q^{\mu_{0,1,2}}$ $=1-(1-p_{d})^{2}e^{-\eta\mu_{0,1,2}},$
with $\eta=\eta_{det}\times 10^{-\frac{\alpha L/2}{10}}$, $\eta_{det}=30\%$ is total detection efficiency, $\alpha=0.2~{}\rm{dB/km}$ is the linear attenuation of a standard optical fibre, and $L$ is the distance.
@@ 2.3.1 Security analysis on TF QKD
For the error rates,
$E^{\mu_{0,1,2}}$ $=\frac{1}{2}+\frac{1-p_{d}}{2Q^{\mu_{0,1,2}}}\left[e^{-\mu_{0,1,2}\eta(1-e_{opt}-E_{m})}-e^{-\mu_{0,1,2}\eta(e_{opt}+E_{m})}\right],$ (5)
$E_{m}$ $=\frac{1}{2}-\frac{\sin(2\pi/M)}{4\pi/M},$
@@
with $e_{opt}=3\%$ is the channel optical error rate. In the presence of the wavelength-switching attack, Alice and Bob estimate $e_{1,U}$, $Y_{0,L}$, and $Y_{1,L}$ using Eq. 3 and Eq. 4, but now calculate $E^{\mu_{0,1,2}}$ and $Q^{\mu_{0,1,2}}$ using the experimentally observed parameters $\mu^{\prime}_{0,1,2}=g\times\mu_{0,1,2}$. We conduct a simulation analysis using parameters corresponding to a typical experiment [22] to illustrate the security threat posed by a wavelength-switching attack on the TF QKD system. The results with and without an attack on the TF QKD system are depicted in Fig. 5. The black solid line represents the lower bound $R_{L}$ as defined by Eq. 1 under no attack conditions. Subsequently, we simulate the degradation of security bounds due to Eve’s wavelength-switching attack. Specifically, in Fig. 5, the red-dashed and blue-dotted lines respectively indicate values of incorrectly estimated ($R_{L}$) by Alice and Bob when $g=1.087$ in the presence of an attack. As shown in Fig. 5 compared to no-attack scenario represented by black solid line, the secure key rate ($R_{L}$) given by red-dashed line is significantly higher. In other words, Alice and Bob overestimate their secure key rate in response to an attack. The gap between these two curves indicates that Eve can eavesdrop on a portion of secure key rate without being detected while eventually deducing shared secure key between Alice and Bob.
@@
In Fig. 6, we present the impact of the wavelength-switching factor $f_{\Delta}$ on the lower bound of secure key rate at a fixed distance ($560~{}\mathrm{k}\mathrm{m}$) and $g=1.087$. The graph illustrates both incorrect and correct estimates of secure key rates for the TF QKD system under attack, represented by upward triangles and downward triangles respectively. In this scenario, an increase in parameter $f_{\Delta}$ results in a higher mean photon number for quantum states prepared by Alice and Bob. This consistently elevated mean photon number leads to an overestimation of the lower bound $R_{L}$ of key rate, as indicated by the red-square markers, thereby widening the gap between overestimated and correct key rates illustrated by blue-dot markers. To summarize, wavelength-switching attacks cause Alice and Bob to overestimate their key rate, compromising the security of a TF QKD system.
2.3.2 Security analysis on SNS-TF-QKD
@@
Next, we consider the case of SNS-TF-QKD [10]. Similar to the previous example, it is assumed that both Alice and Bob employ three distinct intensities, namely $\mu_{0}$, $\mu_{1}$, and $\mu_{2}$. In the finite-key case, the lower bound for the secure key rate can be derived from Ref. [37, 38, 23].
$R_{L}=\frac{1}{N}\left\{n_{1}\left[1-H_{2}(e_{1}^{ph})\right]-fn_{t}H_{2}(E_{z})\right\}-\gamma,$ (6)
where $N$ represents the finite size, $n_{1}$ denotes the lower bound of the number of survived untagged bits, $e_{1}^{ph}$ gives the upper bound of the phase-flip error rate for those survived untagged bits, $f$ is the error correction inefficiency, $n_{t}$ is the number of survived untagged bits, $E_{z}$ is the corresponding bit-flip error rate in those survived bits, and $\gamma$, which takes logarithmically small values, is for security considerations with finite-data size and advanced decoy state analysis.
To evaluate Eq. 6, Alice and Bob need to calculate the lower bound of the expected value of the counting rate of untagged photons
@@
where $S_{jk}\left(jk=\left\{00,10,01,20,02\right\}\right)$ denotes the counting rate of source $jk$ [37]. In the wavelength-switching attack, the parameters $S_{jk}$ should be calculated based on the experimentally obtainable data $\mu^{\prime}_{0}$, $\mu^{\prime}_{1}$, and $\mu^{\prime}_{2}$. According to Ref. [37, 23], $n_{1}$ and $n_{t}$ can be obtained from $S_{1,L}$, $E_{z}$ can be calculated based on $S_{jk}$, and $e_{1}^{ph}$ can be estimated from $e_{1,U}$. The upper bound of the expected value of the phase-flip error rate satisfies [39]
$e_{1,U}$ $\leq\frac{T_{\Delta}-0.5e^{-2\mu_{1}}S_{00}}{2\mu_{1}e^{-2\mu_{1}}S_{1,L}},$ (8)
where $T_{\Delta}$ is the effective rate of detectors. Furthermore, we can obtain a much lower QBER $E_{z}$ in those survived bits by applying active odd-parity pairing [40, 41, 42] to the above method.
Figure 7: The secure key rate as a function of distance for SNS-TF-QKD is observed to be $1$ and $1.087$ for two different values of the multiplicative factor $g$. The simulation parameters utilized in the calculations are as follows. $N=5.4\times 10^{14}$, $P_{d}=1.4\times 10^{-11}$, $f=1.16$, $\alpha=0.157$, and $\eta_{det}=0.6$.
@@ 3.1 Countermeasures
The observed variation in the intensity of quantum states is attributed to the change in optical insertion loss. This change is mainly because the AOM receives an uncharacterized frequency modulation and insufficient driving voltage when the FLC adjusts it to match the reference light’s wavelength switching as analyzed in Section 2.2.1. Being immune to the wavelength-switching attack, we propose a calibration procedure for the AOM before system operation to minimize the insertion loss due to driving voltage. In step 1, the driving voltage of the AOM is scanned under a fixed modulation frequency to determine the optimal voltage value, $V_{o}$, corresponding to the minimum insertion loss at this certain modulation frequency. In step 2, the modulation frequency of the AOM is scanned with $V_{o}$ found in Step 1, obtaining the optimal frequency $F_{o}$, corresponding to the further minimized insertion loss. Afterwards, $V_{o}$ and $F_{o}$ are set as the basic parameters for the FLC to control the AOM. This process ensures that the wavelength-switching attack will only increase the insertion loss of the OPLL, thus mitigating the associated security risks.
An additional countermeasure could be involving monitoring devices to detect eavesdropping. For example, during heterodyne measurements, photodetectors PD1 and PD2 might be used to identify multiple peaks in the spectrum, potentially indicating a wavelength-switching attack. Moreover, by monitoring the intensity with PD5 and PD6 in the source stations, the system can monitor the intensity fluctuation of Alice’s and Bob’s output light. These fluctuations can be indicative of the potential wavelength-switching attack. From another aspect of view, inspired by the fully passive QKD protocol [43, 44], the vulnerability related to wavelength-switching attacks might be mitigated if the wavelength/optical field of a laser source can be passively modulated or locked.
@@
3.2 Conclusion
The TF-class QKD has been demonstrated to surpass the fundamental rate-distance limit, which necessitates the generation of identical optical fields by two independent lasers located at separate positions for Alice and Bob. In current implementations, a third-party reference light is commonly employed to lock the wavelengths of the lasers in order to achieve matching. Although this reference-light-locking structure does not directly contribute to key generation, it introduces a potential vulnerability that can be exploited by an eavesdropper, posing a security risk to the devices used by Alice and Bob. This study unveils a wavelength-switching attack on the TF-class QKD system employing an optical OPLL structure. Experimental results demonstrate that Eve can exploit this attack to reduce attenuation in the AOM of the TF-class QKD system, causing optical power from Alice and Bob to exceed intended levels. Importantly, this attack challenges fundamental assumptions in TF-class QKD protocols regarding mean photon number of prepared quantum states and well-protected source stations. Despite altering wavelengths of lights emitted by Alice and Bob, these hacked source units remain consistent, enabling stable interference at Charlie’s location. Theoretical analysis conducted on both TF QKD protocol and SNS-TF-QKD protocol indicates that the wavelength-switching attack could significantly compromise security in implementation of TF-class QKD - legitimate users overestimate secure key rates when facing such attacks. Since this attack is due to imperfect calibration in the external-reference-locking configuration, we propose countermeasures via loss-minimization calibration or using no-external-reference-light scheme [45, 46, 47, 48, 35]. The possibility of this attack on various OPLLs schems shall be experimentally verified in the future. Additionally, the practical security of other variants of TF-QKD, such as phase-matching QKD [9], shall be investigated as future study.
4 Methods
4.1 Model of AOM